What are the center and radius of the circle whose equation is x2 + (y-3)2 = 144?
Answer by Cromlix(4381)

x^2 + (y - 3)^2 = 144
Multiply out the bracket.
x^2 +(y^2 - 6y + 9) = 144
x^2 + y^2 - 6y + 9 - 144 = 0
x^2 + y^2 -6y - 135 = 0
Centre (0, 3) radius =sqrrt 0^2 + 3^2 + 135
= sqrrt 144
= 12
Hope this helps.
:-)